1. You are working for Dunder Mifflin, a company in Scranton, PA that manufactures paper.  Corporate has decided that they need to increase sales by 20%.  Suppose you know that the own price elasticity of demand for paper in your market is ‐0.5.  What pricing strategy do you suggest?  

A.  Increase the price of paper by 40%.

B.  Decrease the quantity of advertising by 40%.

C.  Decrease the price of paper by 40%.

D.  Decrease the price of paper by 20%.

E.  Shut down the Scranton branch.

2.  Old McDonald has a farm.  On his farm he has some sheep.  He also has some cows.  He makes only two goods: wool from the sheep and milk from the cows.  He has a typically shaped production possibilities frontier.  He recently took out a loan and purchased a fancy automated sheep shearing machine which doubles his hourly wool production.   What happens to McDonald’s marginal opportunity cost of milk (in terms of wool) as a result of the new machine?

A.  It stays the same.

B.  It decreases.

C.  It increases.

D.  We cannot tell from the above information.

E.  It necessarily becomes linear.

.3  Consider the following scenario regarding the gasoline market in the U.S.  A crisis in the Middle East shuts down some oil production.  At the same time, people in the U.S. start buying a lot more fuel efficient cars.  Based on just this information, what do we know for sure would happen in the gasoline market?

A.  The equilibrium price would decrease.

B.  The equilibrium quantity would decrease.

C.  The equilibrium price would increase.

D.  The equilibrium quantity would increase.

E.  Both equilibrium quantity and equilibrium price would increase

Hannah’s Sporting company makes Badminton rackets.  They are currently producing 24,800 rackets a month using 100 hours of labor and 15 units of capital.  The marginal product of labor at 100 hours is 54 rackets.  The marginal product of capital at 15 units of machinery is 96 rackets.  Hannah’s input prices are as follows.  The price of labor is $9 per hour and the price of capital is $24 per unit of machinery.  Hannah is happy with the level of output, but she wants to be sure she is minimizing cost. She should

A.  do nothing, her costs are minimized already.

B.  use more capital only since it is more productive.

C.  use less labor only since it is not very productive.

D.  use more labor and less capital.

E.  use less labor and more capital.
Answer #1

Question 1) option a)

↑ Total Revenue by a5% (uce el d4 dd--b ed71 cdnelastic demand, IER 11 ubve 0 Rule venue thw puica by 40g Rence TR 个by 4067, dus option A)

Question 2 ) option c).

Since now for producing the same amount of milk, more of wool has to be given up, as efficiency of wool machine has risen, thus opportunity cost of milk production rises.

Question 3) option b)

As both supply & demand of oil falls, thus equilibrium quantity falls

Bcoz now more fuel efficient cars are being used, thus demand for oil falls

Question 4) option D)

At cost minimising combination

Ratio of MP = ratio of prices

So MP_L/ MP_K = P_L/ P_K

= Thus MP_L/P_L = MP_K / P_K

Now MP of Labor divided by it's price equals 54/9 = 6

& For capital it is 96/24 = 4

As MP_L / P_L is greater than MP_K / P_K

So use more Labor & less k